OpenStudy (anonymous):

answer quick!!! The city of baltimore has been declining in population for the last fifty years. in the year 2000, the population of baltimore was 651 thousand and declining at a rate of .75% per year if this trend continues, give a formula for the population of baltimore, p, in thousands, as a function of years, t, since 2000

OpenStudy (anonymous):

Turn your brain on, quick!!!! Let P(t) be the population in years t after 2000. The population in 2000 is 651,000 After one year, the population P(1) = 651,000 ( 1 - 0.75/100) = 651,000 (1 - 0.0075) = 651,000 (0.9925) After two years, the population is P(2) = P(1) ( 1 - 0.75/100) = P(1) (0.9925) = 651,000 (0.9925) (0.9925) = 651,000 (0.9925)^2 So what is P(t) in general? If you can't see it, write out expressions for P(3) and P(4).